rectangles length is 10 feet longer than its width. If the area of the rectangle is 75 square feet, find its demensions
Answer by Greenfinch(383)

Call width x, then length is x + 10
Area is x(x+ 10) = 75
x^2 + 10x - 75 = 0
(x +15)(x - 5) = 0
x = 15 or -5
-5 is not an option, so width is 5 and length is 15